OpenStudy (anonymous):

The hypotenuse of a right triangle is 10 centimeters, and the shorter leg is 5 centimeters. Find the length of the other leg.

OpenStudy (anonymous):

let it b x, then x^2+y^2=h^2 x^2=10^2-5^2 x=root75 =5root3

OpenStudy (anonymous):